# Binary Cumulative Encoding meets Time Series Forecasting

### v2

@@ Abstract
Recent studies in time series forecasting have explored formulating regression via classification task. By discretizing the continuous target space into bins and predicting over a fixed set of classes, these approaches benefit from stable training, robust uncertainty modeling, and compatibility with modern deep learning architectures. However, most existing methods rely on one-hot encoding that ignores the inherent ordinal structure of the underlying values. As a result, they fail to provide information about the relative distance between predicted and true values during training. In this paper, we propose to address this limitation by introducing binary cumulative encoding (BCE), that represents scalar targets into monotonic binary vectors. This encoding implicitly preserves order and magnitude information, allowing the model to learn distance-aware representations while still operating within a classification framework. We propose a convolutional neural network architecture specifically designed for BCE, incorporating residual and dilated convolutions to enable fast and expressive temporal modeling. Through extensive experiments on benchmark forecasting datasets, we show that our approach outperforms widely used methods in both point and probabilistic forecasting, while requiring fewer parameters and enabling faster training.
@@ 1 Introduction
Time series forecasting plays a key role in decision-making across domains such as finance [akita2016deep], healthcare [song2020distributed], retail [bandara2019sales], and climate science [zaini2022systematic]. While classical methods like ARIMA and Exponential Smoothing remain widely used in practice [hyndman2018forecasting], recent advances increasingly favor deep learning approaches [lim2021time, li2024deep], particularly in scenarios requiring high-precision forecasting [zhang2024probts].
@@
A critical component of deep learning-based forecasting is data preprocessing. This typically includes normalization, min-max or mean scaling [salinas2020deepar], Box-Cox transformation [hyndman2018forecasting], and short-time Fourier transforms [furman2023short]. Recent studies [ansari2024chronos, liu2024timer] have explored quantization, discretizing the continuous target space into fixed-size bins. Values are then represented as one-hot vectors, enabling a tokenized view of time series aligned with Transformer-style architectures. Empirical results suggest that such representations enhance performance, especially within Time Series Foundation Models [zhang2024probts].
@@
However, one-hot encoding lacks ordinal awareness, treating bins as unrelated categories and ignoring relative distances–important cues for modeling temporal dynamics. To address this, we propose Binary Cumulative Encoding (BCE): a representation where all bins below (or equal to) the current value are activated. This monotonic encoding, previously applied to tabular data [gorishniy2022embeddings], preserves global ordinal structure and yields richer inputs for neural models. Figure 1 shows an example of BCE applied to a time series.
Figure 1: Visualization of different representations of time series data. (left) Original time series showing monthly airline passenger counts. (center) One-hot encoding of discretized values. (right) Binary Cumulative Encoding (BCE), where all bins less than or equal to the active bin are highlighted, preserving ordinal structure.
@@
Choosing an appropriate neural architecture is another central design decision. RNNs, CNNs, and Transformers are commonly used, with Transformer-based models [zhang2024probts] excelling due to their capacity for long-range temporal modeling [vaswani2017attention]. Yet, their nearly quadratic complexity with sequence length [tay2022efficient] poses practical challenges.
@@
To overcome this, we introduce a CNN-based model that combines 2D and 1D convolutions to process BCE representations efficiently. This design reduces computational cost while preserving representational expressiveness. Moreover, BCE naturally supports probabilistic output by modeling distributions over bins–an ability often missing in standard CNN forecasting models.
We validate our method on the ProbTS benchmark [zhang2024probts], which includes diverse datasets from multiple domains. Our model consistently outperforms strong baselines in both point and probabilistic forecasting tasks.
The main contributions of this work are:
  • A novel discrete representation, Binary Cumulative Encoding (BCE), which captures ordinal and metric structure in time series.
  • A lightweight convolutional model, BinConv, combining 2D and 1D convolutions to process BCE sequences efficiently.
  • Extensive evaluation on the ProbTS benchmark, showing that our method achieves state-of-the-art results across multiple forecasting datasets.
@@ 2 Related Work
Classical and Deep Learning Methods for Time Series Forecasting. Classical models such as ARIMA, exponential smoothing, and vector autoregression are widely used for their simplicity and efficiency [hyndman2018forecasting]. However, they often underperform on large-scale datasets exhibiting complex and nonlinear patterns, as shown in M4, M5, M6 forecasting competitions [makridakis2018m4, makridakis2022m5, makridakis2024m6]. Deep learning approaches improve upon these by framing forecasting as a sequence-to-sequence problem using fixed context windows [lim2021time]. Recent developments like Time Series Foundation Models [liang2024foundation, das2024decoder, darlow2024dam] leverage large-scale pretraining to generalize across domains, achieving strong results in both point and probabilistic forecasting [zhang2024probts].
Neural Network Architectures for Time Series Forecasting. Neural models vary in their trade-offs between accuracy, scalability, and speed. Transformer-based models such as PatchTST [nie2023a], TFT [lim2021temporal], and Informer [zhou2021informer] capture long-range dependencies well, but incur high computational costs due to their quadratic scaling with sequence length. Earlier RNN-based models like LSTM [hochreiter1997long], GRU [che2018recurrent], and DeepAR [salinas2020deepar] are more efficient at inference, though limited by their sequential nature, which hinders training parallelism. CNN-based architectures such as TCN [bai2018empirical], DeepTCN [chen2020probabilistic], and ModernTCN [luo2024moderntcn] offer fast, parallel training and strong forecasting performance. Following this line, our CNN-based BinConv model balances accuracy and efficiency when used with BCE inputs.
Token-Based Representations of Time Series. Inspired by NLP, recent work reformulates time series forecasting as next-token prediction using discretization. Models like LLMTime [gruver2023large], TiMER [liu2024timer], and Chronos [ansari2024chronos] encode values as digits, patches, or bins. Gorishniy et al. [gorishniy2022embeddings] proposed piecewise linear encoding (PLE) for tabular data, preserving ordinal structure. Most models rely on one-hot encoding, which discards distance and order information between bins. This limits the ability to learn value relationships. In contrast, our BCE encoding captures both order and magnitude, improving learning while remaining compatible with token-based models.
Probabilistic Forecasting with Deep Learning Methods. Probabilistic forecasting offers both predictions and uncertainty. Deep learning models approach this via quantile regression (e.g., TFT, DeepAR), parametric output distributions (e.g., DeepTCN), or generative models like TimeGrad [rasul2021autoregressive] and GRU-NVP [rasul2021multivariate]. While effective, generative approaches can be unstable and costly to train. Token-based models such as Chronos and TiMER enable probabilistic inference via sampling in discrete space. Our model adopts this approach using BCE, supporting both point and probabilistic forecasts through structured sampling.
3 Binary Cumulative Encoding
@@
Figure 2: Example of BCE encoding for 4 bins.
The scale of time series can vary significantly, even within a single dataset. Therefore, before converting time series values into an encoding space, we apply normalization to map values into a quantized range. Similar to the Chronos model [ansari2024chronos], we use the mean scaling technique [salinas2020deepar], where each value is divided by the mean over a given context window. Specifically, we compute the scaled value as $\tilde{x}=x/s$, where $s=\frac{1}{C}\sum_{i=1}^{C}\left|x_{i}\right|$.
@@
$\mathrm{BCE}(x)$ $=[v_{1},\dots,v_{D}]\in\mathbb{R}^{D}$
$v_{d}$ =$=\begin{cases}0,&\tilde{x}<b_{d}\\ 1,&\tilde{x}\geq b_{d}\end{cases},$ (1)
where $b_{d}$ indicates a $d$-th bin in the original space. An example of construction of BCE is shown in Figure 2.
The interval from $b_{0}$ to $b_{D}$ is uniformly split into $D$ bins in a data-agnostic manner, where $D$, $b_{0}$, and $b_{D}$ are hyperparameters. In our experiments, we do not fine-tune these values per dataset and use fixed settings of $D=1000$, $b_{0}=-5$, and $b_{D}=5$, unless stated otherwise. This quantization results in binary cumulative encoding of the form $111\ldots 100\ldots 0$.
This type of encoding is not new in the context of tabular deep learning for numerical features [gorishniy2022embeddings]. However, unlike previous work, we not only transform the inputs using this scheme, but also forecast targets directly in the encoding space, as detailed in Section 4.2. After generating a forecast in the encoding space, we apply an inverse transformation by assigning the average value of the bin encoding, i.e., $\tilde{y}=(b_{i}+b_{i+1})/2$, where $i$ is the index of the last $1$. Finally, to obtain the forecast in the original space, we reverse the mean scaling: $y=s\cdot\tilde{y}$. To the best of our knowledge, this approach is novel in the context of time series forecasting.
@@ 4.1 Architecture
Input $\mathbb{R}^{C}$ BCE $\mathbb{R}^{C\times D}$ $M$ BinConv block $\mathbb{R}^{C\times D}$ Conv1d Sigmoid $\mathbb{R}^{D}$ Inverse BCE $\mathbb{R}$
@@
Figure 3: BinConv architecture. (a) First, we preprocess the time series using BCE. Then, the BinConv model is trained in the BCE space, and during forecasting, the inverse BCE transformation is applied to obtain predictions in the original space. BinConv consists of $M$ stacked BinConv blocks followed by a final 1D convolution layer with a large kernel size. (b) Each BinConv block is the core component of the architecture, comprising three convolutional layers and a residual connection. The DyTanh activation function is used as a substitute for layer normalization. The output dimension of each BinConv block matches its input dimension.
@@
Conv2d $(C,\;s_{1})$ $1\ /\ K$ $1$
Conv1d-1 $s_{2}$ $K\ /\ K$ $K$
Conv1d-2 $s_{2}$ $K\ /\ C$ $K$
[…]
To model time series effectively in the discrete BCE space, the architecture must preserve ordinal relationships, support autoregressive forecasting, and maintain computational efficiency. To this end, we propose BinConv, a convolutional neural architecture that efficiently captures the local temporal and monotonic structure of BCE vectors by combining 2D and 1D convolutions.
The BinConv model consists of $M$ convolutional blocks followed by a final convolutional layer with a sigmoid activation function, producing an output vector in the BCE space (see Figure 3(a)). For each convolution, we add padding on both sides to preserve output dimensionality. To preserve the cumulative structure of the BCE representation, we use ones for left padding and zeros for right padding. Notably, we do not use fully connected layers in the model. In fact, including fully connected layers degrades performance, as further discussed in Section 6.1.
Each BinConv block comprises three convolutional layers (see Figure 3(b)). The first is a 2D convolution with kernel size $(C,3)$, where $C$ corresponds to the context length. This layer transforms the input matrix–obtained after applying BCE–into $K$ one-dimensional vector representations, where $K$ is the number of output channels and treated as a hyperparameter. Following the work of [zhu2025transformers], we use the DyTanh activation function, which replaces layer normalization in our architecture.
Next, we apply two 1D convolutional layers with ReLU activation, followed by dropout after the second layer. Inspired by TCN architectures [bai2018empirical, luo2024moderntcn], we employ depthwise convolution with the number of groups set to $K$, which significantly reduces computational cost without sacrificing performance 6.2. Each BinConv block is designed such that the input and output dimensions are identical. This allows the use of a residual connection as the final operation in the block, enabling the stacking of multiple BinConv blocks while mitigating the vanishing gradient problem. The main parameters of the model are listed in Table 1.
To train the model, we apply the binary cross-entropy loss to the model output. This approach is analogous to a multi-label classification setup, where each class is predicted independently using a "one-versus-all" scheme.
@@ 4.2 Forecasting
To forecast time series, we apply an autoregressive approach: forecasting at time step $t+1$, appending the result to the context, then forecasting at $t+2$, and so on. To do this, we need to transform the vector of sigmoid-activated outputs into a BCE. We interpret each component of the output as the success probability of an independent Bernoulli trial. Due to the design of the BCE, only sequences of the form $11\ldots 10\ldots 0$ are considered valid. Therefore, we normalize probability of each valid sequence by the total probability of all valid sequences, denoted as $Z$.
More formally, the probability of a valid encoding vector $e_{m}$, where the first $m$ components are 1 and the remaining $D-m$ components are 0, is given by:
@@
000
$\hat{p}(e_{0})=0.048$
$p(e_{0})=\frac{0.048}{Z}\approx 0.109$
100
$\hat{p}(e_{1})=0.032$
$p(e_{1})=\frac{0.032}{Z}\approx 0.073$
110
$\hat{p}(e_{2})\approx 0.288$
$p(e_{2})=\frac{0.288}{Z}\approx\mathbf{0.654}$
111
$\hat{p}(e_{3})=0.072$
$p(e_{3})=\frac{0.072}{Z}\approx 0.164$
$Z=\sum_{m}\hat{p}(e_{m})=0.048+0.032+0.288+0.072=0.44$
Figure 4: Converting the neural network output $[0.4,0.9,0.2]$ into a BCE, the most probable sequence is $[1,1,0]$.
@@ 5.1 Setup
In this paper, our primary focus is on univariate forecasting. We evaluate performance on three univariate datasets: M4 Daily, M4 Weekly, and Tourism Monthly [makridakis2020m4, athanasopoulos2011tourism]. The prediction horizon $H$ for each dataset is set to $14$, $13$, and $24$, respectively. The context length is set to $C=3H$ for all datasets.
We adopt the training and evaluation setup from [zhang2024probts]. Specifically, we train each model for $50$ epochs, then select the best checkpoint based on performance on the training set, and report evaluation metrics on the test set. For all models, we used the Adam optimizer [kingma2015adam] for training. We compute CRPS using $20$ quantile levels and $100$ samples to assess the quality of probabilistic forecasts, and NMAE to evaluate point forecasting performance following to [zhang2024probts]. All computations were done on NVIDIA Tesla A10 22.5GB GPU RAM.

### v3

# BinConv: A Neural Architecture for Ordinal Encoding in Time-Series Forecasting Andrei Chernov1, Vitaliy Pozdnyakov2, Ilya Makarov3 1

@@ Abstract
Recent work in time series forecasting has explored reformulating regression as a classification task. By discretizing the continuous target space into bins and predicting over a fixed set of classes, these approaches benefit from more stable training, improved uncertainty modeling, and compatibility with modern deep learning architectures. However, most existing methods rely on one-hot encoding, which ignores the inherent ordinal structure of the target values. As a result, they fail to convey information about the relative distance between predicted and true values during training. In this paper, we address this limitation by applying Cumulative Binary Encoding (CBE), a monotonic binary representation that transforms both model inputs and outputs. CBE implicitly preserves ordinal and magnitude information, allowing models to learn distance aware representations while operating within a classification framework. To leverage CBE effectively, we propose BinConv, a fully convolutional neural network architecture designed for probabilistic forecasting. We demonstrate that standard fully connected layers are not only less computationally efficient than convolutional layers when used with CBE, but also degrade forecasting performance. Our experiments on standard benchmark datasets show that BinConv achieves superior performance compared to widely used baselines in both point and probabilistic forecasting, while requiring fewer parameters and enabling faster training.
@@ 1 Introduction
Time series forecasting plays a key role in decision-making across domains such as finance (Akita et al. 2016), healthcare (Song et al. 2020), retail (Bandara et al. 2019), and climate science (Zaini et al. 2022). While classical methods like ARIMA and Exponential Smoothing remain widely used in practice (Hyndman and Athanasopoulos 2021), recent advances increasingly favor deep learning approaches (Lim and Zohren 2021; Li and Law 2024), particularly in scenarios requiring high-precision forecasting (Zhang et al. 2024).
@@
A critical component of deep learning-based forecasting is data preprocessing. This typically includes normalization, min-max or mean scaling (Salinas et al. 2020), Box-Cox transformation (Hyndman and Athanasopoulos 2021), and short-time Fourier transforms (Furman et al. 2023). Recent studies (Ansari et al. 2024; Liu et al. 2024) have explored discretizing the continuous target space into fixed-size bins. Values are then represented as one-hot vectors, enabling a tokenized view of time series aligned with Transformer-style architectures. Empirical results suggest that such representations enhance performance, especially within Time Series Foundation Models (Zhang et al. 2024).
@@
However, one-hot encoding lacks ordinal awareness, treating bins as unrelated categories and ignoring relative distances, which are important for modeling temporal dynamics. To address this, we propose Cumulative Binary Encoding (CBE): a representation where all bins below (or equal to) the current value are activated. This monotonic encoding, previously applied to tabular data (Gorishniy, Rubachev, and Babenko 2022), preserves global ordinal structure and yields richer inputs for neural models.
@@
Choosing an appropriate neural architecture is another central design decision. RNNs, CNNs, and Transformers are commonly used, with Transformer based models (Zhang et al. 2024) excelling due to their capacity for long range temporal modeling (Vaswani et al. 2017). Yet, their nearly quadratic complexity with sequence length (Tay et al. 2022) poses practical challenges.
@@
As a result, in this paper we introduce BinConv, a CNN-based model that combines two dimensional and one dimensional convolutions to process CBE representations efficiently. Moreover, BinConv with CBE naturally supports probabilistic output by modeling distributions over bins, a capability often missing in standard CNN forecasting models. Remarkably, CBE and BinConv are effective only when used together. Our ablation study in Section 6 shows that neither CBE combined with a Transformer model nor one-hot encoding used with BinConv achieves satisfactory performance.
@@ 2 Related Work
Neural Network Architectures for Time Series Forecasting. Neural models vary in their trade-offs between accuracy, scalability, and speed. Transformer-based models such as PatchTST (Nie et al. 2023), TFT (Lim et al. 2021), and Informer (Zhou et al. 2021) capture long-range dependencies well, but incur high computational costs due to their quadratic scaling with sequence length. Earlier RNN-based models like LSTM (Hochreiter and Schmidhuber 1997), GRU (Che et al. 2018), and DeepAR (Salinas et al. 2020) are more efficient at inference, though limited by their sequential nature, which hinders training parallelism. CNN-based architectures such as TCN (Bai, Kolter, and Koltun 2018), DeepTCN (Chen et al. 2020), and ModernTCN (Luo and Wang 2024) offer fast, parallel training and strong forecasting performance. Following this line, our CNN-based BinConv model balances accuracy and efficiency when used with CBE inputs.
Token-Based Representations of Time Series. Inspired by NLP, recent work reformulates time series forecasting as next-token prediction using discretization. Models like LLMTime (Gruver et al. 2023), TiMER (Liu et al. 2024), and Chronos (Ansari et al. 2024) encode values as digits, patches, or bins. Gorishniy et al. (Gorishniy, Rubachev, and Babenko 2022) proposed piecewise linear encoding (PLE) for tabular data, preserving ordinal structure. Most models rely on one-hot encoding, which discards distance and order information between bins. This limits the ability to learn value relationships. In contrast, our CBE captures both order and magnitude, improving learning while remaining compatible with token-based models.
3 Cumulative Binary Encoding
@@
Figure 1: Example of CBE encoding for 4 bins.
The scale of time series can vary significantly, even within a single dataset. Therefore, before converting time series values into an encoding space, we apply normalization to map values into a quantized range. Similar to the Chronos model (Ansari et al. 2024), we use the mean scaling technique (Salinas et al. 2020), where each value is divided by the mean over a given context window. Specifically, we compute the scaled value as $\tilde{x}=x/s$, where $s=\frac{1}{C}\sum_{i=1}^{C}\left|x_{i}\right|$.
@@
$\mathrm{CBE}(x)$ $=[v_{1},\dots,v_{D}]\in\mathbb{R}^{D}$
$v_{d}$ $=\begin{cases}0,&\tilde{x}<b_{d}\\ 1,&\tilde{x}\geq b_{d}\end{cases},$ (1)
where $b_{d}$ indicates a $d$-th bin in the original space. An example of construction of CBE is shown in Figure 1.
The interval from $b_{0}$ to $b_{D}$ is uniformly split into $D$ bins in a data-agnostic manner, where $D$, $b_{0}$, and $b_{D}$ are hyperparameters. In our experiments, we do not fine-tune these values per dataset and use fixed settings of $D=1000$, $b_{0}=-5$, and $b_{D}=5$, unless stated otherwise. This quantization results in Cumulative Binary Encoding of the form $111\ldots 100\ldots 0$.
This type of encoding is not new in the context of tabular deep learning for numerical features and can be viewed as a simplified version of the piecewise linear encoding proposed in (Gorishniy, Rubachev, and Babenko 2022). However, unlike previous work, we not only transform the inputs using this scheme, but also forecast targets directly in the encoding space, as detailed in Section 4.2. After generating a forecast in the encoding space, we apply an inverse transformation by assigning the average value of the bin encoding, i.e., $\tilde{y}=(b_{i}+b_{i+1})/2$, where $i$ is the index of the last $1$. Finally, to obtain the forecast in the original space, we reverse the mean scaling: $y=s\cdot\tilde{y}$. To the best of our knowledge, this approach is novel in the context of time series forecasting.
@@ 4.1 Architecture
Input $\mathbb{R}^{C}$ CBE $\mathbb{R}^{C\times D}$ $M$ × BinConv block $\mathbb{R}^{C\times D}$ Conv1d Sigmoid $\mathbb{R}^{D}$ Inverse CBE $\mathbb{R}$
@@
Figure 2: BinConv architecture. (a) First, we preprocess the time series using CBE. Then, the BinConv model is trained in the CBE space, and during forecasting, the inverse CBE transformation is applied to obtain predictions in the original space. BinConv consists of $M$ stacked BinConv blocks followed by a final 1D convolution layer with a large kernel size. (b) Each BinConv block is the core component of the architecture, comprising three convolutional layers and a residual connection. The DyTanh activation function is used as a substitute for layer normalization. The output dimension of each BinConv block matches its input dimension.
@@
Conv2d $(C,\;s_{1})$ 1/K1\ /\ K 11
Conv1d-1 $s_{2}$ K/KK\ /\ K $K$
Conv1d-2 $s_{2}$ K/CK\ /\ C $K$
Conv1d $s_{3}$ C/ 1C\ /\ 1 11
To model time series effectively in the discrete CBE space, the architecture must preserve ordinal relationships, support autoregressive forecasting, and maintain computational efficiency. To this end, we propose BinConv, a convolutional neural architecture that efficiently captures the local temporal and monotonic structure of CBE vectors by combining 2D and 1D convolutions.
The BinConv model consists of $M$ convolutional blocks followed by a final convolutional layer with a sigmoid activation function, producing an output vector in the CBE space (see Figure 2(a)). For each convolution, we add padding on both sides to preserve output dimensionality. To preserve the cumulative structure of the CBE representation, we use ones for left padding and zeros for right padding.
Each BinConv block comprises three convolutional layers (see Figure 2(b)). The first is a 2D convolution with kernel size $(C,3)$, where $C$ corresponds to the context length. This layer transforms the input matrix–obtained after applying CBE–into $K$ one-dimensional vector representations, where $K$ is the number of output channels and treated as a hyperparameter. Following the work of (Zhu et al. 2025), we use the DyTanh activation function, which replaces layer normalization in our architecture.
Next, we apply two 1D convolutional layers with ReLU activation, followed by dropout after the second layer. Inspired by (Liu et al. 2022; Luo and Wang 2024), we employ depthwise convolution which is a special case of grouped convolution where the number of groups equals the number of channels $K$. This significantly reduces computational cost without sacrificing performance (see Section 6.4). Each BinConv block is designed such that the input and output dimensions are identical. This allows the use of a residual connection as the final operation in the block, enabling the stacking of multiple BinConv blocks while mitigating the vanishing gradient problem. The main parameters of the model are listed in Table 1.
To train the model, we apply the binary cross-entropy loss to the model output. This approach is analogous to a multi-label classification setup, where each class is predicted independently using a ”one-versus-all” scheme.
@@ 4.2 Forecasting
To forecast time series, we apply an autoregressive approach: forecasting at time step $t+1$, appending the result to the context, then forecasting at $t+2$, and so on. To do this, we need to transform the vector of sigmoid-activated outputs into a CBE. We interpret each component of the output as the success probability of an independent Bernoulli trial. Due to the design of the CBE, only sequences of the form $11\ldots 10\ldots 0$ are considered valid. Therefore, we normalize probability of each valid sequence by the total probability of all valid sequences, denoted as $Z$.
More formally, the probability of a valid encoding vector $e_{m}$, where the first $m$ components are 1 and the remaining DmD-m components are 0, is given by:
@@
000 $\hat{p}(e_{0})=0.048$ $p(e_{0})=\frac{0.048}{Z}\approx 0.109$ 100 $\hat{p}(e_{1})=0.032$ $p(e_{1})=\frac{0.032}{Z}\approx 0.073$ 110 $\hat{p}(e_{2})\approx 0.288$ $p(e_{2})=\frac{0.288}{Z}\approx\mathbf{0.654}$ 111 $\hat{p}(e_{3})=0.072$ $p(e_{3})=\frac{0.072}{Z}\approx 0.164$ $Z=\sum_{m}\hat{p}(e_{m})=0.048+0.032+0.288+0.072=0.44$
Figure 3: Converting the neural network output $[0.4,0.9,0.2]$ into a CBE, the most probable sequence is $[1,1,0]$.
@@ 5.1 Setup
We evaluate performance on three univariate datasets: M4 Daily (daily), M4 Weekly (weekly), and Tourism Monthly (tourism) (Makridakis, Spiliotis, and Assimakopoulos 2020; Athanasopoulos et al. 2011). The prediction horizon $H$ for each dataset is set to $14$, $13$, and $24$, respectively. The context length is set to $C=3H$ for all datasets.
Additionally, we report long-term forecasting results on several multivariate datasets commonly used in the research community: ETTh1, ETTh2, ETTm1, ETTm2, and Illness (Zhang et al. 2024). The context length $C$ and prediction horizon $H$ were set to $C=H=36$ for the Illness dataset and $C=H=96$ for the others.
For the baseline models, we used the optimal hyperparameters provided in (Zhang et al. 2024). In contrast, BinConv was evaluated using its default configuration without any fine-tuning. This configuration is consistent across univariate and multivariate datasets, with the exception of the number of bins (see Table 2). For multivariate datasets, we reduced the number of bins to accelerate experiment runtime.
We adopt the training and evaluation setup from (Zhang et al. 2024). Specifically, we train each model for $50$ epochs, and report evaluation metrics on the test set. For all models, we used the Adam optimizer (Kingma 2015) for training.
We evaluate point forecasts using the normalized mean absolute error (NMAE) and probabilistic forecasts using the continuous ranked probability score (CRPS), both widely adopted in time series forecasting research (Zhang et al. 2024; Chen et al. 2020; Ansari et al. 2024; Rasul et al. 2021b). The NMAE is defined as
where $K$ is the dimensionality of the time series ($K=1$ in the univariate case) and $T$ is the prediction horizon. The CRPS is given by
$\operatorname{CRPS}(F^{-1},x)=\int_{0}^{1}2\Lambda_{\alpha}\left(F^{-1}(\alpha),x\right)\,d\alpha,$
with quantile loss $\Lambda_{\alpha}(q,z)=(\alpha-\mathbb{I}\{z<q\})(z-q)$ and indicator function $\mathbb{I}\{\cdot\}$. We compute CRPS using $19$ quantile levels and $100$ samples to assess the quality of probabilistic forecasts, and NMAE to evaluate point forecasting performance following to (Zhang et al. 2024).
For BinConv, as discussed in Section 3, we applied mean scaling independently to each sample prior to quantization. This was necessary because certain datasets, such as the M4 benchmark, consist of time series with widely varying scales.
All computations were performed on NVIDIA Tesla A10 GPU with 22.5 GB of RAM using CentOS 7 operating system, and all models were implemented using PyTorch 2.7.
5.2 Baseline Models
We select four representative baselines. First, we include PatchTST (Nie et al. 2023), a transformer based model that achieves the best average performance across datasets in the ProbTS benchmark (Table 10 in (Zhang et al. 2024)). Second, we evaluate DLinear (Zeng et al. 2023), an efficient model in terms of both runtime and parameter count, while still achieving competitive performance.
Unlike BinConv, PatchTST and DLinear are non-autoregressive models: they predict the entire forecast horizon in a single inference pass. This design enables faster inference but lacks flexibility, as each change in forecast horizon requires retraining. In contrast, autoregressive models can adapt to arbitrary horizons at inference time. Additionally, PatchTST and DLinear are not probabilistic models and do not provide uncertainty estimates.
To address this gap in baselines, we include two autoregressive probabilistic models: GRU-NVP (Rasul et al. 2021b), which combines a GRU encoder with normalizing flows; and TimeGrad (Rasul et al. 2021a), which applies score based diffusion modeling to generate future trajectories by denoising noise conditioned on past observations.
For all four baselines, we adapt the hyperparameter configurations provided by the ProbTS benchmark.
We exclude foundation time series models from our baselines, as their zero shot performance remains inferior to that of state-of-the-art local models (Zhang et al. 2024; Toner et al. 2025).